OpenStudy (anonymous):

What is the mean of 22, 17, 60, 54, 31, 92, 23? Round to the nearest tenth of a unit. 42.7 31.7 54.7 60.7

OpenStudy (anonymous):

will medal

OpenStudy (anonymous):

@Mike7888 @marylou004

OpenStudy (igreen):

The mean is the average.. add up all the terms and divide by 7(the amount of terms there are).

OpenStudy (anonymous):

uhhhhhhhhh

OpenStudy (igreen):

\(\dfrac{22+ 17+ 60+ 54+ 31+ 92+ 23}{7} =~?\)

OpenStudy (igreen):

Add all the top numbers, and then divide by 7. \(\dfrac{22+ 17+ 60+ 54+ 31+ 92+ 23}{7}\) 22+ 17+ 60+ 54+ 31+ 92+ 23 = 299 \(\dfrac{299}{7}\) Can you divide 299 / 7? @iamawesome1

OpenStudy (anonymous):

42.7 is you answer aim

OpenStudy (igreen):

@Mike7888 Don't give out answers.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

42.71428571428571

OpenStudy (igreen):

Yep, that rounds to 42.7, which is A.

OpenStudy (anonymous):

good job :)

OpenStudy (anonymous):

every one here deserves medals

OpenStudy (igreen):

Including you..lol.

OpenStudy (anonymous):

mike i gave @iGreen a medal i think he gave you one

OpenStudy (anonymous):

yep u right @iGreen

OpenStudy (igreen):

Yep, I did.

OpenStudy (anonymous):

ok thanks i will hand some out

OpenStudy (anonymous):

so mike7888 can give who ever he chosses a medal to

OpenStudy (anonymous):

i gave iam one...cause you went and got the whole answer lol

OpenStudy (anonymous):

we all got medals now i got 4 smart cents how many do you have @iGreen and @Mike7888

OpenStudy (anonymous):

10 lol

OpenStudy (igreen):

I have 997.

OpenStudy (anonymous):

i cn't see how many smart cents every one has cause this computer i am on is unable to

OpenStudy (igreen):

And I've spent 550.

OpenStudy (anonymous):

damn @iGreen

OpenStudy (igreen):

Lol. I'm a 98..so yeah. :P

OpenStudy (anonymous):

I'm working on it

OpenStudy (anonymous):

how do i use my smart cents

OpenStudy (anonymous):

lets move on to the next question every one